# No evidence for $H_{0}$ anisotropy from Tully–Fisher or supernova distances

### v1

@@ 2.3 Peculiar velocity data
In C15, the luminosity-weighted density field is reconstructed from the redshift-space positions of galaxies in the 2M++ catalogue using the iterative method of Yahil et al. (1991). 2M++ is a whole-sky redshift compilation of $69160$ galaxies (Lavaux & Hudson, 2011), derived from 2MASS photometry (Skrutskie et al., 2006) and redshifts from 2MRS (Huchra et al., 2012), 6dF (Jones et al., 2009), and Sloan Digital Sky Survey DR7 (Abazajian et al., 2009). Apparent magnitudes are corrected for Galactic extinction, $k$-corrections, evolution, and surface brightness dimming. The catalogue is magnitude-limited to $K<11.5$ in the 2MRS region and $K<12.5$ in the 6dF and Sloan Digital Sky Survey regions.
@@ 3 Methodology
In particular our method jointly calibrates the Tully–Fisher relation relation (or supernovae) with the galaxy bias and the peculiar velocity field, following the methodology developed in S25. While B25 also largely adopts a forward-modelling approach, one key step does not fall within that framework: by querying peculiar velocity in redshift space, they use the observed redshift to assign peculiar velocity to a galaxy, approximately updating the fiducial velocity field calibration of C15, and then use that to estimate the cosmological redshift. That is a backward-modelling step (going from the redshift observation to a model parameter, the cosmological redshift or analogously distance) that introduces complications in triple-valued zones where a single observed redshift can map to multiple line-of-sight distances (Strauss & Willick, 1995). We instead model the observed redshift directly by querying the peculiar velocity field in real space. Here, we provide a brief summary of the model; a more detailed explanation is given in Appendix˜A, with further discussion in S25.
@@ 3.1 Tully–Fisher model
The distances and true linewidths of each galaxy are latent model parameters over which we marginalise. The true linewidths are drawn from an inferred Gaussian distribution (Bartlett & Desmond, 2023) and compared to the observations through a likelihood function. Via the Tully–Fisher relation, the true linewidths determine the absolute magnitudes, which, combined with the distances, yield predictions of the apparent magnitudes to be compared with the data. Likewise, the true model distances, together with the peculiar velocity model, provide predictions of the observed redshifts for comparison with the measurements. The forward model accounts for both homogeneous and inhomogeneous Malmquist bias. The homogeneous case is an assumption that sources are uniformly distributed in volume, yielding a distance prior $p(r)\propto r^{2}$ in the absence of selection effects. The inhomogeneous Malmquist bias is set by the density field of C15 and modelled as
$b(r)=1+b_{1}\delta(r),$ (5)
where $\delta(r)$ is the density contrast at the galaxy’s position. To ensure non-negative values, $b(r)$ is clipped from below at zero. We adopt this linear bias model to be consistent with the reconstruction of C15, which uses linear theory to relate the galaxy density field, effectively smoothed over scales of $4~h^{-1}\,\mathrm{Mpc}$, to a peculiar velocity field. We treat $b_{1}$ as a model parameter and infer it with a wide uniform prior.
In general, Tully–Fisher relation samples are subject to a complex selection function combining HI flux, linewidth, optical magnitude, and possibly redshift selection. Following the empirical approach of Lavaux (2016), we model the prior distribution of galaxy distances as
$\pi(r\mid\bm{\theta})=\frac{b(r,\,\bm{\theta})\,f(r,\,\bm{\theta})}{\int\mathrm{d}r^{\prime}\,b(r^{\prime},\,\bm{\theta})\,f(r^{\prime},\,\bm{\theta})},$ (6)
where $b(r,\,\bm{\theta})$ accounts for the inhomogeneous Malmquist bias through the large-scale density field, $\bm{\theta}$ collectively denotes the model parameters, and
$f(r,\,\bm{\theta})=r^{p}\exp\!\left[-\left(\frac{r}{R}\right)^{n}\right].$ (7)
Here $p$, $n$, and $R$ are free parameters: $p\approx 2$ recovers the homogeneous Malmquist bias, $R$ sets the characteristic scale of sample incompleteness, and $n$ controls how sharply the completeness falls off. The normalisation in Eq.˜6 is computed explicitly since it depends on $\bm{\theta}$. This treatment remains phenomenological, approximating the selection rather than modelling it directly. A rigorous forward modelling of the survey selection, which requires knowledge of the selection function, is presented in Kelly et al. (2008) and we recently applied it to $H_{0}$ inference in Stiskalek et al. (2025b).
@@
where $\bm{\Delta}_{\mathrm{ZP}}$ is the zero-point dipole vector and $\hat{\bm{r}}$ is the unit vector in the direction of each galaxy. The dipole amplitude is denoted $\Delta_{\rm ZP}$, with direction specified in Galactic coordinates as $(\ell_{\Delta_{\rm ZP}},\,b_{\Delta_{\rm ZP}})$. The zero-point $a_{\rm TFR}$ is perfectly degenerate with the Hubble constant $H_{0}$, since the distance modulus depends on the logarithm of the luminosity distance. This leads to a degenerate parameter combination $a_{\rm TFR}+5\log h$, where $h\equiv H_{0}/(100~\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1})$. Following S25 and similar works, we express distances in units of $h^{-1}\,\mathrm{Mpc}$, rendering the analysis independent of $h$ and constraining only the degenerate combination.
As explored by B25, one may allow for a dipole in this degenerate parameter across the sky. Provided sufficient sky coverage, such a dipole can be constrained. This introduces four possible interpretations:
@@
    Spurious $H_{0}$ anisotropy arising from flows not captured by C15.
B25 found no evidence for significant spatial variation in linewidths within the CosmicFlows-4 W1 subsample, by comparing the linewidth distribution of sources in the northern and southern hemispheres of the ALFALFA survey (Haynes et al., 2018). Moreover, the use of WISE photometry minimises systematic variation in magnitude calibration across the sky. Based on these considerations, they interpreted their measured dipole in $a_{\rm TFR}+5\log h$ as a potential signature of anisotropy in $H_{0}$. Assuming the dipole in $a_{\rm TFR}+5\log h$ arises from a dipole in $H_{0}$, the corresponding fractional variation in the Hubble constant is given by
@@ 3.2 Pantheon+ supernova model
On the other hand, when using the covariance matrix of Pantheon+ supernovae from Lane et al. (2024), we explicitly sample the Tripp parameters $\mathcal{A}$ and $\mathcal{B}$ to predict the supernova apparent magnitude (see Eq. 3), since their data does not assume a fiducial Tripp standardisation or its contribution to the covariance matrix. We evaluate the likelihood in the apparent magnitudes as in Eq. 5 of Seifert et al. (2025), except that we explicitly sample the galaxy distances, rather than setting them deterministically from the observed redshift which is equivalent to assuming no redshift uncertainty and no peculiar velocities.
@@ 3.3 Galactic extinction
A potential systematic effect in inferring the dipole in the zero-point, or equivalently in the apparent magnitude, arises from the treatment of Galactic extinction. We test this on the CosmicFlows-4 subsample, for which the applied Galactic extinction corrections are available in the public data release. The reported magnitudes in a given waveband $\lambda$ are given by (see Section 2.4 of Kourkchi et al. 2019):
@@
$A_{k}^{\lambda}$ is the $k$-correction, $A_{a}^{\lambda}$ is the total flux aperture correction, and E(BV)E(B\!-\!V) is the colour excess (reddening). As described by Kourkchi et al. (2019), the CosmicFlows-4 catalogue uses the Schlegel et al. (1998) $100~\mu\mathrm{m}$ cirrus maps to extract the Milky Way E(BV)E(B\!-\!V) values. For wavebands $\lambda\in(u,g,r,i)$, the extinction coefficients $R_{\lambda}$ are taken from Schlafly & Finkbeiner (2011), while for the infrared bands, $R_{\rm W1}=0.186$ and $R_{\rm W2}=0.123$, as determined by Fitzpatrick (1999).
To assess potential systematics from Galactic extinction corrections, we use the dustmaps package [^1] https://dustmaps.readthedocs.io/en/latest/ (Green, 2018) to extract E(BV)E(B\!-\!V) values at the angular positions of CosmicFlows-4 galaxies. We adopt the extinction maps of Chiang (2023) and Planck Collaboration et al. (2016). The former is a dust reddening map on the plane of the sky, derived from a reanalysis of Schlafly & Finkbeiner (2011), which in turn is based on Schlegel et al. 1998. It uses tomographically constrained templates from Wide-field Infrared Survey Explorer galaxy density fields to remove contamination from the cosmic infrared background (CIB). The latter map, from Planck Collaboration et al. (2016), employs the generalized needlet internal linear combination (GNILC) method to separate Galactic dust emission from CIB anisotropies, yielding an alternative all-sky extinction estimate. We consider these approaches for modelling Galactic extinction:
  • Use the E(BV)E(B\!-\!V) values from Schlegel et al. (1998), jointly sampling the extinction coefficient $R_{\rm W1}$. We adopt a Gaussian prior on $R_{\rm W1}$ centred at 0.19, with either a standard deviation of 0.01, consistent with the measurement of Yuan et al. 2013, or a broader, more conservative prior with standard deviation 0.05.
  • Use the E(BV)E(B\!-\!V) values from Chiang (2023) or Planck Collaboration et al. (2016), also jointly sampling $R_{\rm W1}$ under the same prior choices as above.
@@ 3.4 Mock data generation
For $N$ sources, we draw sky positions uniformly over the sphere, excluding a mock “Zone of Avoidance” defined by $|b|<7.5^{\circ}$. Each line of sight is then used to query the C15 field for the density and peculiar velocity. The radial distance $r$ is sampled from the prior in Eq.˜6, which incorporates the linear galaxy bias of Eq.˜5. At this distance we evaluate the line-of-sight peculiar velocity $\bm{v}(\bm{r})$ and compute the true redshift,
@@ 3.5 Inference procedure
To sample the posterior distribution we use the No-U-Turn Sampler (No-U-Turn Sampler; Hoffman & Gelman 2011) method of Hamiltonian Monte Carlo as implemented in the numpyro [^2] https://num.pyro.ai/en/latest/ package (Phan et al., 2019; Bingham et al., 2019), ensuring a Gelman–Rubin statistic $\hat{R}-1\leq 0.001$ for convergence (Gelman & Rubin, 1992). We adopt a uniform prior on the zero-point dipole amplitude from 0 to 0.3 and sample its direction uniformly over the surface of a sphere. For all remaining parameters, we adopt uniform priors over sufficiently large ranges, except for $\sigma_{v}$, $\sigma_{\rm int}$ for which we adopt reference (and scale invariant) priors $\pi(x)\propto 1/x$. All these parameters are shared between both the isotropic and dipole models so their choice does not affect the Bayes factors between the models.
@@
where $\bm{\theta}$ is the set of free parameters of the model, $\mathcal{L}$ is the likelihood function and $\pi(\bm{\theta})$ is the prior on the parameters. We estimate this using the harmonic package [^3] https://github.com/astro-informatics/harmonic, which fits a normalising flow to the posterior samples to compute evidence using the harmonic estimator (McEwen et al., 2021; Polanska et al., 2024). For the Tully–Fisher relation samples we numerically marginalise over both galaxy latent parameters ($r$ and $\eta_{\rm true}$) at each Markov Chain Monte Carlo step, yielding a low-dimensional posterior suitable for evidence computation with harmonic. In contrast, for the Pantheon+ samples the covariance prevents straightforward numerical marginalisation of the distances, requiring them to be sampled explicitly. This results in an approximately $600$-dimensional posterior, which makes evidence calculation challenging. We therefore restrict evidence computation to the Tully–Fisher relation samples. For these, we compare models using the Bayes factor, defined as the ratio of evidences for two models $M_{1}$ and $M_{2}$,
@@
$\ell_{\Delta a_{\rm TFR}}$ $127^{\circ}$ Zero-point dipole Galactic longitude
$b_{\Delta a_{\rm TFR}}$ $10^{\circ}$ Zero-point dipole Galactic latitude
@@
$\beta^{\star}$ $0.43$ Velocity scaling of C15
@@
$n$ $1.4$ Distance prior incompleteness rate
Table 1: Injected parameter values used to generate mock Tully–Fisher relation catalogues mimicking the CosmicFlows-4 W1-band sample, assuming the C15 density and peculiar velocity field. The parameters correspond to the posterior mean from the dipole inference on the observed sample. A comparison of the resulting redshift distribution for one mock catalogue with the CosmicFlows-4 subsample is shown in Fig.˜1.

### v2

@@ 2.3 Peculiar velocity data
In Carrick_2015, the luminosity-weighted density field is reconstructed from the redshift-space positions of galaxies in the 2M++ catalogue using the iterative method of Yahil_1991. 2M++ is a whole-sky redshift compilation of $69160$ galaxies (Lavaux_2011), derived from 2MASS photometry (Skrutskie_2006) and redshifts from 2MRS (Huchra_2012), 6dF (Jones_2009), and Sloan Digital Sky Survey DR7 (Abazajian_2009). Apparent magnitudes are corrected for Galactic extinction, $k$-corrections, evolution, and surface brightness dimming. The catalogue is magnitude-limited to $K<11.5$ in the 2MRS region and $K<12.5$ in the 6dF and Sloan Digital Sky Survey regions.
@@ 3 Methodology
In particular our method jointly calibrates the Tully–Fisher relation relation (or supernovae) with the galaxy bias and the peculiar velocity field, following the methodology developed in VF_olympics. While Boubel_2025 also largely adopts a forward-modelling approach, one key step does not fall within that framework: by querying peculiar velocity in redshift space, they use the observed redshift to assign peculiar velocity to a galaxy, approximately updating the fiducial velocity field calibration of Carrick_2015, and then use that to estimate the cosmological redshift. That is a backward-modelling step (going from the redshift observation to a model parameter, the cosmological redshift or analogously distance) that introduces complications in triple-valued zones where a single observed redshift can map to multiple line-of-sight distances (StraussWillick_1995). We instead model the observed redshift directly by querying the peculiar velocity field in real space. Here, we provide a brief summary of the model; a more detailed explanation is given in Appendix˜A, with further discussion in VF_olympics.
@@ 3.1 Tully–Fisher model
The distances and true linewidths of each galaxy are latent model parameters over which we marginalise. The true linewidths are drawn from an inferred Gaussian distribution (MNR) and compared to the observations through a likelihood function. Via the Tully–Fisher relation, the true linewidths determine the absolute magnitudes, which, combined with the distances, yield predictions of the apparent magnitudes to be compared with the data. Likewise, the true model distances, together with the peculiar velocity model, provide predictions of the observed redshifts for comparison with the measurements. The forward model accounts for both homogeneous and inhomogeneous Malmquist bias. The homogeneous case is an assumption that sources are uniformly distributed in volume, yielding a distance prior $p(r)\propto r^{2}$ in the absence of selection effects. The inhomogeneous Malmquist bias is set by the density field of Carrick_2015 and modelled as
$n(r)=1+b_{1}\delta(r),$ (5)
where $\delta(r)$ is the density contrast at the galaxy’s position. To ensure non-negative values, $n(r)$ is clipped from below at zero. We adopt this linear bias model to be consistent with the reconstruction of Carrick_2015, which uses linear theory to relate the galaxy density field, effectively smoothed over scales of $4~h^{-1}\,\mathrm{Mpc}$, to a peculiar velocity field. We treat $b_{1}$ as a model parameter and infer it with a wide uniform prior.
In general, Tully–Fisher relation samples are subject to a complex selection function combining HI flux, linewidth, optical magnitude, and possibly redshift selection. Following the empirical approach of Lavaux_Virbius, we model the prior distribution of galaxy distances as
$\pi(r\mid\bm{\theta})=\frac{n(r,\,\bm{\theta})\,f(r,\,\bm{\theta})}{\int\mathrm{d}r^{\prime}\,n(r^{\prime},\,\bm{\theta})\,f(r^{\prime},\,\bm{\theta})},$ (6)
where $n(r,\,\bm{\theta})$ accounts for the inhomogeneous Malmquist bias through the large-scale density field, $\bm{\theta}$ collectively denotes the model parameters, and
$f(r,\,\bm{\theta})=r^{p}\exp\!\left[-\left(\frac{r}{R}\right)^{q}\right].$ (7)
Here $p$, $q$, and $R$ are free parameters: $p\approx 2$ recovers the homogeneous Malmquist bias, $R$ sets the characteristic scale of sample incompleteness, and $q$ controls how sharply the completeness falls off. The normalisation in Eq.˜6 is computed explicitly since it depends on $\bm{\theta}$. This treatment remains phenomenological, approximating the selection rather than modelling it directly. A rigorous forward modelling of the survey selection, which requires knowledge of the selection function, is presented in Kelly_2008 and we recently applied it to $H_{0}$ inference in CH0.
@@
where $\bm{\Delta}_{\mathrm{ZP}}$ is the zero-point dipole vector and $\hat{\bm{r}}$ is the unit vector in the direction of each galaxy. The dipole amplitude is denoted $\Delta_{\rm ZP}$, with direction specified in Galactic coordinates as $(\ell_{\Delta_{\rm ZP}},\,b_{\Delta_{\rm ZP}})$. The zero-point $a_{\rm TFR}$ is perfectly degenerate with the Hubble constant $H_{0}$, since the distance modulus depends on the logarithm of the luminosity distance. This leads to a degenerate parameter combination $a_{\rm TFR}+5\log h$, where $h\equiv H_{0}/(100~\mathrm{km}\,\mathrm{s}^{-1}\,\mathrm{Mpc}^{-1})$. Following VF_olympics and similar works, we express distances in units of $h^{-1}\,\mathrm{Mpc}$, rendering the analysis independent of $h$ and constraining only the degenerate combination.
As explored by Boubel_2025, one may allow for a dipole in this degenerate parameter across the sky. Provided sufficient sky coverage, such a dipole can be constrained. This introduces four possible interpretations:
@@
    Spurious $H_{0}$ anisotropy arising from flows not captured by Carrick_2015.
Boubel_2025 found no evidence for significant spatial variation in linewidths within the CosmicFlows-4 W1 subsample, by comparing the linewidth distribution of sources in the northern and southern hemispheres of the ALFALFA survey (Haynes_2018). Moreover, the use of WISE photometry minimises systematic variation in magnitude calibration across the sky. Based on these considerations, they interpreted their measured dipole in $a_{\rm TFR}+5\log h$ as a potential signature of anisotropy in $H_{0}$. Assuming the dipole in $a_{\rm TFR}+5\log h$ arises from a dipole in $H_{0}$, the corresponding fractional variation in the Hubble constant is given by
@@ 3.2 Pantheon+ supernova model
On the other hand, when using the covariance matrix of Pantheon+ supernovae from Lane_2024, we explicitly sample the Tripp parameters $\mathcal{A}$ and $\mathcal{B}$ to predict the supernova apparent magnitude (see Eq. 3), since their data does not assume a fiducial Tripp standardisation or its contribution to the covariance matrix. We evaluate the likelihood in the apparent magnitudes as in Eq. 5 of Seifert_2025, except that we explicitly sample the galaxy distances, rather than setting them deterministically from the observed redshift which is equivalent to assuming no redshift uncertainty and no peculiar velocities.
@@ 3.3 Galactic extinction
A potential systematic effect in inferring the dipole in the zero-point, or equivalently in the apparent magnitude, arises from the treatment of Galactic extinction. We test this on the CosmicFlows-4 subsample, for which the applied Galactic extinction corrections are available in the public data release. The reported magnitudes in a given waveband $\lambda$ are given by (see Section 2.4 of Kourkchi_2019):
@@
$A_{k}^{\lambda}$ is the $k$-correction, $A_{a}^{\lambda}$ is the total flux aperture correction, and E(BV)E(B\!-\!V) is the colour excess (reddening). As described by Kourkchi_2019, the CosmicFlows-4 catalogue uses the Schlegel_1998 $100~\mu\mathrm{m}$ cirrus maps to extract the Milky Way E(BV)E(B\!-\!V) values. For wavebands $\lambda\in(u,g,r,i)$, the extinction coefficients $R_{\lambda}$ are taken from Schlafly_2011, while for the infrared bands, $R_{\rm W1}=0.186$ and $R_{\rm W2}=0.123$, as determined by Fitzpatrick_1999.
To assess potential systematics from Galactic extinction corrections, we use the dustmaps package [^1] https://dustmaps.readthedocs.io/en/latest/ (dustmaps) to extract E(BV)E(B\!-\!V) values at the angular positions of CosmicFlows-4 galaxies. We adopt the extinction maps of Chiang_2023 and Planck_2016. The former is a dust reddening map on the plane of the sky, derived from a reanalysis of Schlafly_2011, which in turn is based on Schlegel_1998. It uses tomographically constrained templates from Wide-field Infrared Survey Explorer galaxy density fields to remove contamination from the cosmic infrared background (CIB). The latter map, from Planck_2016, employs the generalized needlet internal linear combination (GNILC) method to separate Galactic dust emission from CIB anisotropies, yielding an alternative all-sky extinction estimate. We consider these approaches for modelling Galactic extinction:
  • Use the E(BV)E(B\!-\!V) values from Schlegel_1998, jointly sampling the extinction coefficient $R_{\rm W1}$. We adopt a Gaussian prior on $R_{\rm W1}$ centred at 0.19, with either a standard deviation of 0.01, consistent with the measurement of Yuan_2013, or a broader, more conservative prior with standard deviation 0.05.
  • Use the E(BV)E(B\!-\!V) values from Chiang_2023 or Planck_2016, also jointly sampling $R_{\rm W1}$ under the same prior choices as above.
@@ 3.4 Mock data generation
For $N$ sources, we draw sky positions uniformly over the sphere, excluding a mock “Zone of Avoidance” defined by $|b|<7.5^{\circ}$. Each line of sight is then used to query the Carrick_2015 field for the density and peculiar velocity. The radial distance $r$ is sampled from the prior in Eq.˜6, which incorporates the linear galaxy bias of Eq.˜5. At this distance we evaluate the line-of-sight peculiar velocity $\bm{v}(\bm{r})$ and compute the true redshift,
@@ 3.5 Inference procedure
To sample the posterior distribution we use the No-U-Turn Sampler (No-U-Turn Sampler; Hoffman_2011) method of Hamiltonian Monte Carlo as implemented in the numpyro [^2] https://num.pyro.ai/en/latest/ package (Phan_2019; Bingham_2019), ensuring a Gelman–Rubin statistic $\hat{R}-1\leq 0.001$ for convergence (Gelman_1992). For each model, we run four independent chains with $1000$ warm-up steps and $2000$ sampling steps, typically yielding more than $4000$ effective samples for the dipole and other parameters. We adopt a uniform prior on the zero-point dipole amplitude from 0 to 0.3 and sample its direction uniformly over the surface of a sphere. For all remaining parameters, we adopt uniform priors over sufficiently large ranges, except for $\sigma_{v}$, $\sigma_{\rm int}$ for which we adopt reference (and scale invariant) priors $\pi(x)\propto 1/x$. All these parameters are shared between both the isotropic and dipole models so their choice does not affect the Bayes factors between the models.
@@
where $\bm{\theta}$ is the set of free parameters of the model, $\mathcal{L}$ is the likelihood function and $\pi(\bm{\theta})$ is the prior on the parameters. We estimate this using the harmonic package [^3] https://github.com/astro-informatics/harmonic, which fits a normalising flow to the posterior samples to compute evidence using the harmonic estimator (McEwan_2021; Polanska_2024). For the Tully–Fisher relation samples we numerically marginalise over both galaxy latent parameters ($r$ and $\eta_{\rm true}$) at each Markov Chain Monte Carlo step, yielding a low-dimensional posterior suitable for evidence computation with harmonic. In contrast, for the Pantheon+ samples the covariance prevents straightforward numerical marginalisation of the distances, requiring them to be sampled explicitly. This results in an approximately $600$-dimensional posterior, which makes evidence calculation challenging. While harmonic is designed to compute the Bayesian evidence directly from posterior samples of moderately high dimension, it has typically been tested on $\mathcal{O}(10)$-dimensional problems. For example, Piras_2024 applied harmonic to a 39-dimensional posterior, validating their results against nested sampling, and also to a 159-dimensional posterior for which nested sampling was computationally infeasible. harmonic uses a normalising flow within a harmonic-mean framework to perform density estimation for evidence computation, but for a $\sim 600$-dimensional posterior the required number of samples becomes prohibitive, convergence cannot be guaranteed, and its use has not been previously explored in such high dimensionality. Accordingly, we do not apply harmonic in this high-dimensional setting. We therefore restrict evidence computation to the Tully–Fisher relation samples. For these, we compare models using the Bayes factor, defined as the ratio of evidences for two models $M_{1}$ and $M_{2}$,
@@
$\ell_{\Delta_{\rm ZP}}$ $127^{\circ}$ Zero-point dipole Galactic longitude
$b_{\Delta_{\rm ZP}}$ $10^{\circ}$ Zero-point dipole Galactic latitude
@@
$\beta^{\star}$ $0.43$ Velocity scaling of Carrick_2015
@@
$q$ $1.4$ Distance prior incompleteness rate
Table 1: Injected parameter values used to generate mock Tully–Fisher relation catalogues mimicking the CosmicFlows-4 W1-band sample, assuming the Carrick_2015 density and peculiar velocity field. The parameters correspond to the posterior mean from the dipole inference on the observed sample. A comparison of the resulting redshift distribution for one mock catalogue with the CosmicFlows-4 subsample is shown in Fig.˜1.